References

[1] Paul Erdös and Daniel J. Kleitman, On coloring graphs to maximize the proportion of multicolored k-edges, Journal of Combinatorial Theory 5 (1968), 164–169.

[2] Noga Alon, C. Kenneth Fan, Daniel J. Kleitman and Jozsef Losonczy, Acyclic matchings, Advances in Mathematics 122 (1996), 234–236.

[3] C.K. Fan and R.M. Green, Monomials and Temperley–Lieb algebras, Journal of Algebra 190 (1997), 498–517.

[4] R.M. Green and R.J. Marsh, Quantized symmetric powers, Quarterly Journal of Mathematics 48 (1997), 479–492.